Glencoe Medical Clinic operates a cardiology care unit and a maternity care unit. Colby Hepworth, the clinic’s administrator, is investigating the charges assigned to cardiology patients. Currently, all cardiology patients are charged the same rate per patient day for daily care services. Daily care services are broadly defined as occupancy, feeding, and nursing care. A recent study, however, revealed several interesting outcomes. First, the demands patients place on daily care services vary with the severity of the case being treated. Second, the occupancy activity is a combination of two activities: lodging and use of monitoring equipment. Since some patients require more monitoring than others, these activities should be separated. Third, the daily rate should reflect the difference in demands resulting from differences in patient type. Separating the occupancy activity into two separate activities also required the determination of the cost of each activity. Determining the costs of the monitoring activity was fairly easy because its costs were directly traceable. Lodging costs, however, are shared by two activities: lodging cardiology patients and lodging maternity care patients. The total lodging costs for the two activities were $5,511,000 per year and consisted of such items as building depreciation, building maintenance, and building utilities. The cardiology floor and the maternity floor each occupy 19,200 square feet. Hepworth determined that lodging costs would be assigned to each unit based on square feet.

To compute a daily rate that reflected the difference in demands, patients were placed in three categories according to illness severity, and the following annual data were collected:

Activity Cost of Activity Activity Driver Quantity
Lodging $ 681,300    Patient days 22,710
Monitoring 2,112,600    Monitoring hours used 30,180
Feeding 476,910    Patient days 22,710
Nursing care 4,143,900    Nursing hours 218,100
Total 7,414,710

The demands associated with patient severity are also provided:

Severity Patient Days Monitoring Hours Nursing Hours
High 7,780 15,080 131,000
Medium 11,370 12,400 73,200
Low 3,560 2,700 13,900

Required:

1. Suppose that the costs of daily care are assigned using only patient days as the activity driver (which is also the measure of output). Compute the daily rate using this unit-based approach of cost assignment. If required, round your answer to the nearest cent.
$ 326.50 per day

2. Compute activity rates using the given activity drivers (combine activities with the same driver). If required, round your answers to the nearest cent.

Rate 1 $ 51 per patient day
Rate 2 $ 70 per monitoring hour
Rate 3 $ 19 per hour of nursing care

3. Compute the charge per patient day for each patient type using the activity rates from Requirement 2 and the demands on each activity. Round your interim calculations and final answers to the nearest cent.

High severity $ per patient day
Medium severity $ per patient day
Low severity $ per patient day
Answer #1

Charge per patient day =(Activity rate*Activity consumed)/Patient days

High severity = (51*7780 + 70*15080 + 19*131000)/7780 = $506.6041 per patient day

Medium Severity = (51*11370 + 70*12400 + 19*73200)/11370 = $249.6631 per patient day

Low Severity = (51*3560+ 70*2700+ 19*13900)/3560 = $178.2753 per patient day